find any publicy traded stock company and find its how large its equity per share price and IPO. how large that publicy traded company equity per share before iPO. we need calculation in excel with formulas.
Answer #1

For large companies like Goldman Sachs the pre IPO total equity per share is total equity available to shareholders divided by total outstanding shares.

Pre IPO Equity per share = 5,35,47,990 / 48,00,000 = 11 dollars approximately

Post IPO Equity per share = 3,43,66,778/ 17,00,000 = 20 dollars approximately